OpenStudy (curry):

confusion with limits lim of (sin(x))/(x) as x approaches 0 = 1 but y is there no limit for lim of (1/(x^2)) as x approaches 0?

OpenStudy (anonymous):

since the reason is that if we put 0 in the limit directly the numerator gets to 0 in sinx/x but in the latter it will turn into 1/0 form which is undetermined........

OpenStudy (curry):

ye but sin of z equals 1

OpenStudy (curry):

wait nvm

OpenStudy (anonymous):

didnt get u???

OpenStudy (curry):

ye but if the denominator is 0 then shouldn't sinx /x also be undefined?

OpenStudy (anonymous):

first in case of sin we have numerator to be taken first na?

OpenStudy (curry):

ye so ?

OpenStudy (anonymous):

try this if u can have patience to see... :) :) :) this is the proof for your doubt about sinx/x actually its a proved hypothesis... but fails in case of 1/x^2

OpenStudy (curry):

so here is my question

OpenStudy (curry):

u can write limts as x is approaching some asymptote right?

OpenStudy (curry):

if that is true then y is that finction along with these functions also undefined lim of ((4+x)^2-16)/x as x approaches 0

OpenStudy (curry):

or lim of abs(x)/x as x approaches 0

OpenStudy (curry):

all three of these functions are undefined

OpenStudy (curry):

i understand y the abs(x)/x is undefined but

OpenStudy (curry):

the 4+x)^2+16/x is a hole

OpenStudy (curry):

so then it still has a continous limit just a hole at that point

OpenStudy (curry):

same with 1/x^2 so x=0 is a limt

OpenStudy (curry):

i mean asymptote

OpenStudy (anonymous):

there are some proofs on these dude... see actually limits are governed by some rules and formulae you have to think how to apply them, not to think how they have come.... I cant tell you rather than that because, no one would ever think off

OpenStudy (curry):

so there should be a asymptotic limit

OpenStudy (anonymous):

yes.............

OpenStudy (curry):

so y isnt there?

OpenStudy (curry):

it should matter if the denominator is 0 and undefined becasue a limit only states that as x approaches 0 not when x is zero

OpenStudy (curry):

it shouldn't *

OpenStudy (dumbcow):

@Curry , yes that is hole so its value is indeterminate or 0/0 but the limit as you approach the hole may have a real value...in fact it is 8 lets look at the graph of sin(x)/x notice its approaching 1 at x=0

OpenStudy (curry):

so this function has a limit even though it's denominator is 0. lim of (x^2-3x+2)/(x^2-4)

OpenStudy (curry):

so the book says there is no limit for the hole

OpenStudy (curry):

i understand the sinx / x

OpenStudy (curry):

y don't the other functions have one

OpenStudy (curry):

but the last function i typed has a limit so how is that possible even though the denominator for that funciton would also be 0 but for the other functions i typed the denominator 0 but limit is undefined??

OpenStudy (dumbcow):

that last function is 0/0 when x=2 , this means there could be a limit when its n/0 then limit is undefined or you could say there is an asymptote

OpenStudy (dumbcow):

if you factor it you get (x-2)(x-1) / (x-2)(x+2) = (x-1)/(x+2) at x=2, it equals 1/4 so the limit is 1/4 even though technically there is a hole

OpenStudy (curry):

exactly so the (4+x)^2-16/x also has a hole

OpenStudy (curry):

but y doesn't that have a limt

OpenStudy (dumbcow):

anyway, the point is when you evaluate a limit and you get 0/0 , that just means you don;t know what the limit is...it doesn't mean its undefined or doesn't exist you don;t have enough info

OpenStudy (curry):

but waht is it isn't 0/0 its n/0 then still u can have limit right

OpenStudy (dumbcow):

it does have a limit at y=8 http://www.wolframalpha.com/input/?i=plot+%28%28x%2B4%29^2+-16%29%2Fx+from+-2+to+2

OpenStudy (curry):

but my book says no limit

OpenStudy (dumbcow):

to find the limit algebraically, you would have to change the form of the function so that you don't get 0/0 when you plug in x=0 \[\frac{(x+4)^{2} -16}{x} = \frac{x^{2} +8x}{x} = x+8\]

OpenStudy (curry):

exactly

OpenStudy (curry):

thats what i thought and i put the limit as 8 as x approaches 0

OpenStudy (curry):

but it says thats wrong

OpenStudy (curry):

it says its undefined

OpenStudy (dumbcow):

umm i guess i would disagree with your book...are you sure its talking about the limit or the value technically, that value at x=0 is undefined

OpenStudy (curry):

the direction was find the limit if it exists

OpenStudy (curry):

and the problem was (4+x)^2-16)/x

OpenStudy (curry):

so then the limit for (1/x^2) as x approaches 0 is +infin right or would i be wrong to enter that on a test

OpenStudy (dumbcow):

yes you are correct...limit = pos infinity

OpenStudy (dumbcow):

any number divided by 0 will give you limit of either +-infinity

OpenStudy (curry):

ok so im not wrong right? here was my source my book doesn't have the answers for the evens so i used the online solution here is the link

OpenStudy (curry):

?

OpenStudy (dumbcow):

hmm not sure, looks like maybe they havn't got into infinity or special 0/0 limits yet and they accept it as limit does not exist at this point --i can't see their entire solution or explanation but full correct answers are +infinity and 8

OpenStudy (dumbcow):

this is a calculus course correct?

OpenStudy (curry):

calc C